OpenStudy (anonymous):

Simplify the square root of 2 multiplied by the cube root of 2 how do i work this out @texaschic101

OpenStudy (anonymous):

@sheelo_mughal @phi @aaronq @waterineyes @johnweldon1993

OpenStudy (anonymous):

\[\large \sqrt{2} \times \sqrt[3]{2} = (2)^{\frac{1}{2}} \times (2)^{\frac{1}{3}}\]

OpenStudy (johnweldon1993):

\[\large \sqrt{2} \sqrt[3]{2}\] Remember that we can write this as \[\large 2^{1/2} 2^{1/3}\] And when we multiply this...we just add the exponents \[\large 1/2 + 1/3 = 5/6\] so we have \[\large 2^{5/6}\]

OpenStudy (anonymous):

And you can here use the indices rule that: \[(x)^a \times (x)^b = (x)^{a+b}\]

OpenStudy (anonymous):

Getting @jadee27p ??

OpenStudy (anonymous):

yes still confused though

OpenStudy (anonymous):

about the 1/2

OpenStudy (anonymous):

Square root means 1/2 Like cube root means 1/3 Generally we must represent sqrt(2) as: \[\huge \sqrt[2]{2}\]

OpenStudy (anonymous):

But these square root brackets tell you that you are doing square root only, so we don't write 2 in front of the brackets.. But for 3, 4, 5th roots, you have to write in front..

OpenStudy (anonymous):

Like cube root you will write as: \[\large \sqrt[3]{x}\] Fourth root you will write as: \[\sqrt[4]{x}\] But square root can be written as : \[Either \quad \sqrt{x} \quad or \quad \sqrt[2]{x}\]

OpenStudy (anonymous):

Getting till here??

OpenStudy (johnweldon1993):

Right, just remember the rule \[\Large \sqrt[b]{a^c} = a^{c/b}\] so \[\Large \sqrt[2]{2^1} = 2^{1/2}]\] and \[\Large \sqrt[3]{2^1} = 2^{1/3}\]

OpenStudy (anonymous):

Look for this one: For \(\sqrt[m]{x}\), you can write it as : \((x)^{\frac{1}{m}}\)..